OpenStudy (anonymous):

Which of the following is a solution to tan(2x)=sqrt3 2pi/3 pi/3 pi/6 5pi/6

OpenStudy (anonymous):

well let's solve for 2x:\[\tan (2x)=\sqrt{3}\] what radian angle has a tangent of\[\sqrt{3}\]? well its 60 degrees or pi/3. You should know that from the 30-60-90 triangle definition.  okay so 2x=pi/3 so what does x equal? well it equals pi/3/2 or pi/6 THE ANSWER IS PI/6